# Topologically Stable Hough Transform

### 09 Mar 2026 v1

@@ 1 Introduction
Figure 1: Left: Ten points sampled from two lines. Middle: The classical Hough transform passes to the space of lines in which every input point dualizes into a sinusoidal curve. The blue point on the left, for instance, is tranformed to the blue curve in the middle. The pixels that the blue curve votes for are marked. In that example, there are two pixels (red) that receive five votes. Right: The centers of the high-count pixels are two lines in primal space that correspond to the two drawn lines.
@@ 2 The score function
For $p=(r\cos\Theta,r\sin\Theta)$, $\ell(r,\Theta)$ is the line through $p$ that is perpendicular to the vector $\overrightarrow{p}$. Note that the set of all lines in ${\mathbb{R}}^{2}$ is in bijection with $M$, except that $\ell(r,0)=\ell(-r,\pi)$, as the space of lines is homemorphic to the open Möbius strip. For simplicity, we will first work with $M$ as a planar domain and discuss the twisted glueing to a Möbius strip later.
@@ 4 Computation
Figure 2: Top: the subdivided $\tilde{\mathrm{S}}_{P}$. Bottom: The superlevelsets for threshold $0.8$, $0.7$, and $0.6$, and the nerve of the boxes (the graph structure in blue). Vertices and edges that enter the graph at this threshold are drawn thicker. We observe that for $0.8$, a new connected component is formed in the upper region. For $0.7$, another isolated vertex is formed; moreover, the upper-right box connects with the lower-left box because of the Möbius geometry, hence, the component formed at $0.8$ merges with another one, yielding a persistence of $0.1$ for the component. At $0.6$, the component formed at $0.7$ also gets merged, yielding a persistence of $0.1$ as well. Note that the graph is connected for $0.6$.
@@ 5 A Use-case
We implemented a prototype of our method in Python, available at https://github.com/JRC-ISIA/TopologicalHoughTransformation. Figure 3 shows an exemplary run of our method on a point cloud sampled with noise from three lines (top-left).
@@ 6 Discussion
We have demonstrated that simple tools from computational geometry and topology have the potential to increase the quality of line detection via the Hough transform. Although the experiments presented are limited to line detection, the proposed method generalizes to arbritary shapes by varying the shape parametrization and the corresponding parameter space. We are planning more extensive tests, including the use of different kernel functions, evaluation on the effect of the kernel size, real image data, and the comparison with state-of-art line detection methods. A preliminary evaluation can be found in Appendix C. As a prerequisite, we will first implement a fast version of our prototype – we are optimistic that the combination of adaptive subdivisions and the fact that persistence in dimension 0 can be computed in almost linear time will allow us to apply our method to large image data sets.
@@ Appendix B Lipschitz constants in a box
Note that if $f$ is a sinusoidal curve consisting of all solutions $(r,\Theta)$ of the equation $x_{p}\cos\Theta+y_{p}\sin\Theta=r$, the vertical distance can be computed in constant time by considering the (up to two) local extrema of $f$ wihin the $x$-range of $B$, and the corners of the $B$.
@@
Therefore, the computation of the local Lipschitz constant for $B$ boils down checking whether the vertical distance for $B$ and the curve of $p\in P$ has vertical distance more than $\sigma$ or not. Writing $C$ for the number of curves with smaller vertical distance then $\sigma$, we immediately get with Lemma B.1 a local Lipschitz constant of
@@ Appendix C Experimental evaluation
For the classical method (implemented in OpenCV [2]), the choice of a threshold is highly dependent on the input data (right in Figure˜3), making it difficult, if not impossible, to identify a single threshold that enables the detection of the correct number of lines. With the following experiment, we demonstrate that the proposed method is more likely to produce a threshold that yields a clear and robust separation. For this experiment, we generated $1000$ images, each containing four randomly parameterized lines $(\Theta,r)$ with 18, 17, 16, and 15 sampled points, respectively. Each point was perturbed by an orthogonal displacement drawn uniformly from $\left[-1,1\right)$. Our method is parametrized using the hat kernel $\kappa_{\text{hat}}$, $\sigma=5$, and $\epsilon=5$. We measure the absolute gap between the last included and the first candidate line excluded using persistence, $\Delta_{\text{Pers}}$ (ours), and vote counts, $\Delta_{\text{Vote}}$ (OpenCV), taking the provided order of candidate lines into account. The results are illustrated in Figure˜4. While our method consistently yields a non-zero interval from which a valid threshold can be selected, the vote-based approach exhibits $\Delta_{\text{Vote}}=0$ for $66.5\%$ of the experiments, thereby preventing the selection of a threshold that correctly recovers the true number of lines.
@@
Further, assuming that both methods detect the optimal number of candidate lines in the generated images, Figure˜5 compares the quality of the detected lines using the Euclidean distance in parameter space, and the absolute deviations in $r$ and $\Theta$ to the ground truth. Across all three evaluation metrics, the results consistently demonstrate a performance advantage of our method over the baseline.
@@
margin: MU: I hope I did no confuse $\sigma$ with $\lambda$ here

### 12 Mar 2026 v2

@@ Abstract
keywords:
Computational topology, Hough transform, persistence, quad-tree subdivision, stability
category:
[…]
@@ 1 Introduction
Figure 1: Left: Ten points sampled from two lines. Middle: The classical Hough transform passes to the space of lines in which every input point dualizes into a sinusoidal curve. The blue point on the left, for instance, is transformed to the blue curve in the middle. The pixels that the blue curve votes for are marked. In that example, there are two pixels (red) that receive five votes. Right: The centers of the high-count pixels are two lines in primal space that correspond to the two drawn lines.
@@ 2 The score function
For $p=(r\cos\Theta,r\sin\Theta)$, $\ell(r,\Theta)$ is the line through $p$ that is perpendicular to the vector $\overrightarrow{p}$. Note that the set of all lines in ${\mathbb{R}}^{2}$ is in bijection with $M$, except that $\ell(r,0)=\ell(-r,\pi)$, as the space of lines is homeomorphic to the open Möbius strip. For simplicity, we will first work with $M$ as a planar domain and discuss the twisted glueing to a Möbius strip later.
@@ 4 Computation
Figure 2: Top: the subdivided $\tilde{\mathrm{S}}_{P}$. Bottom: The super-levelsets for threshold $0.8$, $0.7$, and $0.6$, and the nerve of the boxes (the graph structure in blue). Vertices and edges that enter the graph at this threshold are drawn thicker. We observe that for $0.8$, a new connected component is formed in the upper region. For $0.7$, another isolated vertex is formed; moreover, the upper-right box connects with the lower-left box because of the Möbius geometry, hence, the component formed at $0.8$ merges with another one, yielding a persistence of $0.1$ for the component. At $0.6$, the component formed at $0.7$ also gets merged, yielding a persistence of $0.1$ as well. Note that the graph is connected for $0.6$.
@@ 5 A Use-case
[…]
@@ 6 Discussion
We have demonstrated that simple tools from computational geometry and topology have the potential to increase the quality of line detection via the Hough transform. Although the experiments presented are limited to line detection, the proposed method generalizes to arbitrary shapes by varying the shape parametrization and the corresponding parameter space. We are planning more extensive tests, including the use of different kernel functions, evaluation on the effect of the kernel size, real image data, and the comparison with state-of-art line detection methods. A preliminary evaluation can be found in Appendix C. As a prerequisite, we will first implement a fast version of our prototype – we are optimistic that the combination of adaptive subdivisions and the fact that persistence in dimension 0 can be computed in almost linear time will allow us to apply our method to large image data sets.
@@ Appendix B Lipschitz constants in a box
Note that if $f$ is a sinusoidal curve consisting of all solutions $(r,\Theta)$ of the equation $x_{p}\cos\Theta+y_{p}\sin\Theta=r$, the vertical distance can be computed in constant time by considering the (up to two) local extrema of $f$ within the $x$-range of $B$, and the corners of the $B$.
@@
Therefore, the computation of the local Lipschitz constant for $B$ boils down checking whether the vertical distance for $B$ and the curve of $p\in P$ has vertical distance more than $\sigma$ or not. Writing $C$ for the number of curves with smaller vertical distance then $\sigma$, we immediately get with Lemma  B.1 a local Lipschitz constant of
@@ Appendix C Experimental evaluation
For the classical method (implemented in OpenCV [2]), the choice of a threshold is highly dependent on the input data (right in Figure 3), making it difficult, if not impossible, to identify a single threshold that enables the detection of the correct number of lines. With the following experiment, we demonstrate that the proposed method is more likely to produce a threshold that yields a clear and robust separation. For this experiment, we generated $1000$ images, each containing four randomly parameterized lines $(\Theta,r)$ with 18, 17, 16, and 15 sampled points, respectively. Each point was perturbed by an orthogonal displacement drawn uniformly from $\left[-1,1\right)$. Our method is parametrized using the hat kernel $\kappa_{\text{hat}}$, $\sigma=5$, and $\epsilon=5$. We measure the absolute gap between the last included and the first candidate line excluded using persistence, $\Delta_{\text{Pers}}$ (ours), and vote counts, $\Delta_{\text{Vote}}$ (OpenCV), taking the provided order of candidate lines into account. The results are illustrated in Figure 4. While our method consistently yields a non-zero interval from which a valid threshold can be selected, the vote-based approach exhibits $\Delta_{\text{Vote}}=0$ for $66.5\%$ of the experiments, thereby preventing the selection of a threshold that correctly recovers the true number of lines.
@@
Further, assuming that both methods detect the optimal number of candidate lines in the generated images, Figure 5 compares the quality of the detected lines using the Euclidean distance in parameter space, and the absolute deviations in $r$ and $\Theta$ to the ground truth. Across all three evaluation metrics, the results consistently demonstrate a performance advantage of our method over the baseline.